OpenStudy (anonymous):

Medal A snack food is made by mixing 5lb of popcorn that cost $.80 per pound with caramel that costs $2.40 per pound. How much caramel is needed to make a mixture that costs $1.40 per pound?

OpenStudy (anonymous):

You can set the equation this way: Let x = the amount of caramel required. 5%280.80%29+%2B+x%282.40%29+=+%285%2Bx%29%281.40%29 Simplify and solve for x. 4+%2B+2.40%28x%29+=+7+%2B+1.40%28x%29 Subtract 1.40(x) from both sides. 4+%2B+x+=+7 Subtract 4 from both sides.

OpenStudy (anonymous):

i am sorry i cant understand what you wrote can you please explain to me again

OpenStudy (jennyrlz):

Hello ^.^

OpenStudy (anonymous):

5=lbs of popcorn c=lbs of caramel 80=cost per pound of popcorn (in cents) 240=cost per pound of caramel (in cents) 140=cost of both together (in cents) cost(lbs popcorn)+cost(lbs caramel)=cost(lbs popcorn and caramel) 80(5)+240(c)=140(5+c) 400+240c=700+140c Subtract 400 from both sides 240c=300+140c Subtract 140c from both sides 100c=300 Divide both sides by 100 c=3 Answer: 3 lbs of caramel

OpenStudy (anonymous):

thank you soo much :)

OpenStudy (jennyrlz):

Wow I can't understand what he wrote either lol

OpenStudy (anonymous):

sorry my bad its Let x= the amount of caramel required 5*(0.80)+x(2.40)=(5+x)*(1.40) then simplify and solve for x

OpenStudy (anonymous):

much better ha . i found the equation and tried to copy and paste i guess it didnt work

OpenStudy (anonymous):

xD yupss

OpenStudy (anonymous):

how would you solve for x

OpenStudy (anonymous):

@brucebaner @keeks4481 can you please help me

OpenStudy (anonymous):

I already did it.... it will be 3 x=3 So, the answer is 3 lbs of caramel

OpenStudy (anonymous):

he's right ^^

OpenStudy (anonymous):

ok thank you